Into how much water should you dissolve 80.0 g ethylene glycol in order to obtain a solution with a concentration of 4.9 m?
1
Expert's answer
2017-05-25T11:29:10-0400
M((CH2OH)2)=62 g/mol
n((CH2OH)2)=m/M=80/62=1.29 mol

c=n/V

4.9 M = 1.29 mol / x L ---> x=1.29/4.9=0.26 L

V(H2O)=0.26 L = 260 ml